# active distribution network DER distributed energy resource

### v2

@@ II-C Multiple services provision
One of the main contributions of this work is to determine the flexibility capacity for multiple AS simultaneously. Given that all services are provided by the same DERs, meaning the different service compete for the same power flexibility, the objective must account for the value of the different services. To quantify this value, the different service prices are used. This reflects the willingness of system operators to pay and thus the service value for the system. An important feature of this formulation is that it allows to consider the aggregation of flexibility for different services in a non-hierarchical fashion. The multi-service problem (9) is written as an extension of (5). Note that the GCP power for the different services $p_{0}^{s}$ is formulated here for active power services, as these are the services where time-coupling (e.g. due to limited energy storage) is most relevant. The formulation is however applicable for both active and reactive power services.
$\max_{p^{s},\Omega_{p^{s}_{0}}}$ $\quad\sum_{s}J^{s}(\Omega_{p^{s}_{0}})$ (9a)
s.t. $\forall\zeta\in\mathcal{U},\forall s\in\mathcal{S}\quad\sum_{s}(W^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix})\leq Z\zeta+z_{0}$ (9b)
$G^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix}+B^{s}\zeta+b^{s}_{0}=p^{s}_{0}\quad p^{s}_{0}\in\Omega_{p^{s}_{0}}$ (9c)
@@
This objective does not consider the operational costs of DERs. Section II-E details how the DER costs are accounted for and why they are not included in the objective function. In the remainder of this work, we focus on active power services provided at the GCP. We now specify the different types of services that are considered, where an important distinction must be made between symmetric (e.g. FCR) and asymmetric services (e.g. aFRR). All services must be modeled separately as they might have different prices at the same time step, they have different power and energy requirements and their provision might cause different costs. Symmetric services require that for each flexibility bid with a power rating $p_{0s}$, any realization between $p_{0s}$ and $p_{0s}$ must be feasible. For asymmetric services, a power setpoint between 0 and $p^{{}^{\prime}}_{0s}$ or $p^{{}^{\prime\prime}}_{0s}$ and 0 can be requested. For each service, the determined flexibility capacity at the GCP represents the range of power set points that can be realized by the ADN. A robust representation of the exchanged power is therefore required. Additionally, corresponding to standard market practices, the different services for which flexibility is offered, must be available independently. For linear grid and resource models this problem can be written in a linear way by using the superposition principle to sum the power contributions for the different services independently. A separate equality constraint can then be written for each service, linking the service slack power to the disaggregated contributions of the controllable resources. Geometrically, stacking different services can be seen as determining the maximum volume ellipsoids, weighted with the service prices, for which the Minkowski sum lies within the polytope representing the problem’s constraints. Each service is related to the different constraints and the constraint satisfaction depends on all the services independently, reflecting independent selection and activation of the bids. Symmetric services, ranging from $p^{t}_{0s}$ to $p^{t}_{0s}$ at each time step, can be naturally represented as ellipsoids, with some restrictions on the orientation. For asymmetric services, bids can be either for positive or negative power exchange. Any point between zero and the selected bid must be feasible. A natural representation for these services, based on the ellipsoidal capability aggregation, comes in the form of quadrant ellipsoids, where a positive quadrant ellipsoid will be used to represent a positive asymmetric service. Below the used aggregation sets for the different services are explained.
@@
  • For symmetric services, the flexibility is represented as an ellipsoid. Additionally, the ellipsoid must be restricted to be symmetric around all the axes. Without this restriction, the ellipsoids do not lead to valid FCR ranges. Figure 1 illustrates this condition. The red ellipsoid is not symmetric around the axes, meaning that the capacity for upward regulation does not necessarily match the capacity for downward regulation, as illustrated by the red dots. By restricting the ellipsoids to be symmetric around all axes, we obtain ellipsoids of the blue type, for which the blue dots illustrate the matching up/down regulation. With this restriction, $E^{sm}$ becomes diagonal and $e^{sm}$ vanishes. Finally, without loss of generality, the elements of $E^{sm}$ are restricted to be positive as they represent the maximum symmetric power capacity that can be offered at any time step and both positive and negative power exchanges are included through the realization of $\xi$. Symmetric services are then described by (10).
@@
Next to the power flow caused by the offered flexibility, power flows caused by the baseload are included, namely, the power exchanged at the GCP that is not linked to any AS. In this work, we consider three use cases for the base power, referring to the different TSO-DSO coordination approaches. The base power can be uncontrollable, in which case it simply reflects the power flowing at the GCP as a consequence of the uncontrollable prosumption. Alternatively, the base power can be controlled, for example by a centralized DSO who optimizes the expected energy cost of operating the ADN. In the following, this case is referred to as the ’baseload control’. Both these cases are occurrences of the third DSO-TSO coordination scheme, where the DSO validates and transmits the aggregated bids to the TSO. Finally, we also consider the case of ’self-dispatching’. In this case, the base power at different time steps is restricted to an ellipsoid and controllable resources are activated to balance the uncertain prosumption and limit the uncertainty at the GCP. This can be seen as a special case of the fourth coordination scheme, where the DSO is responsible for its own imbalances and only the remaining flexibility can be offered to the TSO. The modeling of baseload power for these cases is detailed here below.
  • In the self-dispatching case, the goal is to minimize the uncertainty on the baseload perceived at the GCP. The baseload power is represented as an ellipsoid using the approach introduced Section II-A. This naturally reduces the uncertainty on the base power exchange as any point within the ellipsoid must be feasible for any realization of the uncertainty, meaning a larger ellipsoid would reserve a larger portion of the DERs flexibility. This leads to the following set of possible baseload powers, as in (6).
@@
    $p^{b}=(B_{1}^{b}D^{{b}^{-1}}(E^{b}\xi+(e^{b}-b(\zeta)))+B_{2}^{b}(K^{b}\xi+\gamma^{b}+\sum_{u}L^{u}\zeta^{u})$
  • For the controllable baseload power, two new variables $p_{0}^{b}$ and $y^{b}$ are introduced. These represent the expected slack power and controllable injections (14). In this case only the expected value $p^{b}_{0}$ is optimized.
@@ II-D Probabilistic Constraint Satisfaction
The available power flexibility in ADNs depends on the stochastic prosumption, as it influences both the network constraints and the capabilities of DERs. Failing to account for these uncertainties leads to an overestimation of the ADN’s power flexibility. Alternatively, accounting for all possible realizations of these uncertain power injections and robustly ensuring the flexibility capacity is available can be overly conservative. Certain grid operators have already recognized this will be an issue when more flexibility is supplied by DERs [31]. To allow DERs with uncertain capabilities to participate in the AS markets, the grid operator allows actors to offer flexibility services as long as the forecast available capacity is available at least 90% of the time [38] (i.e. according to the P90 requirement). To this end, the stochastic prosumption is handled in a probabilistic way in this work. Note that this availability requirement should be interpreted as a joint chance constraint as merely enforcing the flexibility capacity to be present with 90% probability at each time step will lead to more frequent overbidding. As joint chance constraints are generally hard to solve, we determine a robust uncertainty set for the uncertainty drivers that guarantees a certain coverage $1-\epsilon$ of the empirical probability distribution. For uncertainty drivers with gaussian distributions, exact ellipsoidal uncertainty sets can be obtained [39]. Additionally, we consider the computation of ellipsoidal uncertainty sets with a $1-\epsilon$ coverage for the empirical probability distribution of the available data. A comparison of these approaches with robust ellipsoidal sets covering all scenarios is presented in Section III-A. In this work, we assume a single uncertainty driver $\zeta^{u}$ suffices to model the stochastic fluctuations of all individual injections of a certain type. With the ellipsoidal uncertainty sets, the stochastic injections can be formulated as $p^{u}=\bar{p}^{u}+\Delta\zeta^{u}\>\forall u\in\mathcal{U}$, with $\Delta$ a linear mapping from the uncertainty drivers to the stochastic power injections [^4] Consider for example PV plants. The power can be approximated as $P^{cap}*GHI$, where the global horizontal irradiance ($GHI$) uncertain.. Note that ellipsoidal uncertainty sets can be obtained for each time step separately, considering as stochastic variables $\zeta_{t}^{u}$ with $\zeta_{t}^{u}||\leq 1$ at time $t$ as was done in [26]. However, to guarantee a coverage of the joint probability distribution and jointly satisfy the problem constraints, a single ellipsoidal uncertainty set for $\zeta=[\zeta^{u1},\zeta^{u1}...\>\zeta^{u|\mathcal{U}|}]$ with $\zeta||\leq 1$ should be considered for all uncertainty drivers over the full horizon.
II-E Ensuring Cost-Effectiveness
@@
A sufficient condition for the convexity of the problem is that the costs for the DERs must be linear with respect to the allocated power capacity. This is a reasonable assumption as we show in Appendix -A for the example of BESSs. For linear cost functions, the flexibility cost becomes:
@@
Given that any $\xi||\leq 1$ can be selected, the realized benefit is not known apriori. Secondly, optimizing for the best case as an example would lead to the maximization of a norm, making the problem non-convex. Therefore the objective in (21) is selected as a proxy to maximize the flexibility value. Here,the quantity of flexibility offered at each time step is represented by the diagonal elements of $E^{s}$. Denoting the service prices at time steps $t$ as $p^{s}_{t}$, the objective becomes:
$J^{s}=\sum_{t}p^{s}(t)E^{s}(t,t),\quad J=\sum_{s}J^{s}$ (21)
Finally, to avoid biasing the energy available in the baseload case, two constraints are added.
@@
$\sum_{t}(c^{0}(t)p^{b}_{0}(t)-\sum_{r}|c^{r}(t)^{T}|p^{b}(t)|)\geq 0$ (23)
The first one (22) enforces the integral of the difference of baseload power with respect to the expected net load to be zero. The second (23) ensures the power balancing accounts for operational costs, where $r$ indexes the different DERs. Including these cost-effectiveness considerations, the final problem is given by Problem (24). The impact of this controlled baseload on the problem constraints is modeled through (24m) and the influence of the uncontrollable prosumption and corresponding adjustment of the base power is included in (24g). Note that the subscript $u$ refers to the number of uncertainty sets to ensure the formulation remains generic. In the previous section, based on the work from [26], uncertainty sets were defined per timestep. In Section II-D we show the benefits of a different approach. The influence of all the services and the selected baseload control is combined in (24h) and (24i) where $e^{0}$ and $\gamma$ disappear for the baseload case. We introduce auxiliary variables $\alpha^{i}$ for each constraint and $\alpha^{i}_{s}$ to constrain the impact of the different services on each constraint. Reformulating the problem with robust constraints and restricting the aggregated flexibility set to (quadrant) ellipsoids for all services yields the multiservice aggregation problem (24).
$\max_{E^{s},K^{s}}\quad\sum_{s}\sum_{t}p^{s}(t)E^{s}(t,t)$ (24a)
@@
$\sum_{u}w^{0}_{2,i}L^{0}_{u}-\theta_{u,i}\leq\lambda_{i}$ (24g)
@@
$\sum_{t}(c^{0}(t)p^{b}_{0}(t)-\sum_{r}|c^{r}(t)^{T}|p^{b}(t)|)\geq 0$ (24o)
@@ III Results
This section demonstrates the approach for a number of test cases. First results for a the IEEE33 benchmark system are shown. Then the approach is extended to multiple feeders, interconnected to a common primary substation, showing its effectiveness when aggregating the flexibility of realistic distribution systems. Finally the impact of network constraints is discussed. Common inputs for the service prices and the uncertainty sets are used. Input data was obtained from the Réseau de Transports d’Electricité [41] (for the prices), Deutscher Wetterdienst [42] (for the PV production) and historical residential data from [43] for the loads. Figure 2 shows the inputs used for the case studies.
@@
Figure 3: Uncertainty sets from a) a fitted multivariate gaussian distributions, b) robust with respect to all scenarios, c) 1-$\epsilon$ coverage rate with respect to all scenarios.
III-A Comparison of the Ellipsoidal Uncertainty Sets
In Figure 3, we show the boundaries of the ellipsoidal sets obtained in different ways. In all cases, the considered ellipsoidal uncertainty sets offer a coverage of 90%. First, shown in yellow, we show the robust ellipsoids, considering all the possible uncertainty realizations based on the prediction set. Obtaining robust ellipsoidal sets is a well-known problem [44] and consists of determining the minimum volume ellipsoid containing all points in given set. Next, we consider ellipsoids obtained in a probabilistic way, guaranteeing a $90\%$ coverage with respect to the empirical distributions. A first approach to obtain such a set, shown in blue, is based on a multivariate Gaussian distribution, obtained by fitting the prediction scenarios. As the load and PV uncertainties are not truly normally distributed, this leads to an overly conservative uncertainty set. A second approach assumes the set of past observations is representative for the future realizations. This allows to obtain probabilistic uncertainty sets using a heuristic. By computing the mutual distance between all points and selecting the one with the lowest maximum distance to the $1-\epsilon$ points with the smallest distance to this point. There is no guarantee this will yield the ellipsoid with the smallest volume for a certain coverage $1-\epsilon$, but it leads to significant improvements compared to the robust case. This is shown in green in the figure. Representing the ellipsoids is not possible due to the large number of dimensions. Instead we show the range of possible values at all time steps for the different approach, (i.e. the extreme points in all dimensions). The range of prediction scenarios and historical realizations used to obtain the uncertainty sets are also shown.

### v3

@@ II-C Multiple services provision
One of the main contributions of this work is to determine the flexibility capacity for multiple ASs simultaneously. DERs can provide multiple services, meaning the different services compete for the same power flexibility. An important feature of this formulation is that it allows to consider the aggregation of flexibility for different services in a non-hierarchical fashion. The multi-service problem (11) is written as an extension of (5). Note that the GCP power for the different services $p_{0}^{s}$ is formulated here for active power services, which this work focuses on, as these are the services where time-coupling (e.g. due to limited energy storage) is most relevant. The formulation is however applicable for both active and reactive power services. Section II-E details how the objective is chosen to maximise the aggregation set, how DER costs are accounted for and why they are not included in the objective function.
$\max_{p^{s},\Omega_{p^{s}_{0}}}$ $\quad\sum_{s}J^{s}(\Omega_{p^{s}_{0}})\quad\textrm{s.t.}\quad\forall\zeta\in\mathcal{U},\forall s\in\mathcal{S}$ (11a)
$\sum_{s}(W^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix})\leq M_{z}\zeta+z_{0}$ (11b)
$G^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix}+M_{b}^{s}\zeta+b^{s}_{0}=p^{s}_{0},\quad p^{s}_{0}\in\Omega_{p^{s}_{0}}$ (11c)
@@
All services must be modelled separately as they might have different prices at the same time step, they have different power and energy requirements, they may have different time resolutions and activation requirements and their provision might cause different costs. For each service, the determined flexibility capacity at the GCP represents the range of power set points that can be realized by the ADN. Additionally, corresponding to standard market practices, the different services for which flexibility is offered, must be available independently. A robust representation of the exchanged power is therefore required. For linear grid and resource models this problem can be written in a linear way by using the superposition principle to sum the power contributions for the different services independently. A separate equality constraint can then be written for each service, linking the service slack power to the disaggregated contributions of the controllable resources. Geometrically, stacking different services can be seen as determining the maximum volume ellipsoids, weighted with the service prices, for which the Minkowski sum lies within the polytope representing the problem’s constraints. Each service is related to the different constraints and the constraint satisfaction depends on all the services independently, reflecting independent selection and activation of the bids. This superposition also allows to specify additional constraints for individual services. Services with a different time resolution than the base time step can be forced to have the same values in blocks of time steps. DERs with specific characteristics can be excluded from providing certain services, by simply enforcing the power contribution for a specific service to be zero. Considering the different types of services, an important distinction must be made between symmetric (e.g. FCR) and asymmetric services (e.g. aFRR). Symmetric services, ranging from $p^{t}_{0s}$ to $p^{t}_{0s}$ at each time step, can be naturally represented as ellipsoids, with some restrictions on the orientation. For asymmetric services, bids can be either for positive or negative power exchange, and any point between zero and the selected bid $p^{t}_{0s}$ must be feasible. A natural representation for these services, based on the ellipsoidal capability aggregation, comes in the form of quadrant ellipsoids. The representation of these (quadrant) ellipsoids is further detailed here below.
@@
  • For symmetric services, the flexibility is represented as an ellipsoid that is restricted to be symmetric around all the axes. Without this restriction, the ellipsoids do not lead to valid FCR ranges. Figure 1 illustrates this condition. The red ellipsoid is not symmetric around the axes, meaning that the capacity for upward regulation does not necessarily match the capacity for downward regulation, as illustrated by the red dots. By restricting the ellipsoids to be symmetric around all axes, we obtain ellipsoids of the blue type, for which the blue dots illustrate the matching up/down regulation. With this restriction, $E^{sm}$ becomes diagonal and $e^{sm}$ vanishes, leading to (12). Finally, without loss of generality, the elements of $E^{sm}$ are restricted to be positive as they represent the maximum symmetric power capacity that can be offered at any time step and both positive and negative power exchanges are included through the realization of $\xi$.
@@
Next to the offered flexibility, the power exchanged at the GCP that is not linked to any AS, called baseload power, is also included. This work considers three use cases for the base power, referring to the different TSO-DSO coordination approaches. The base power can be uncontrollable, in which case it simply reflects the power flowing at the GCP as a consequence of the uncontrollable prosumption. Alternatively, the base power can be controlled, for example by a centralized DSO who optimizes the expected energy cost of operating the ADN. In the following, this case is referred to as the ’baseload control’. Both these cases are occurrences of the third DSO-TSO coordination scheme, where the DSO validates and transmits the aggregated bids to the TSO. Finally, in the case of ’self-dispatching’ the base power at different time steps is restricted to an predefined (ellipsoidal) set and controllable resources are activated to balance the uncertain prosumption and limit the uncertainty at the GCP. This means the DSO is responsible for its own imbalances and only the remaining flexibility can be offered to the TSO. The modelling of baseload power for these cases is detailed here below.
  • In the self-dispatching case, the goal is to minimize the uncertainty on the baseload perceived at the GCP. The baseload power is represented as an ellipsoid using the approach introduced Section II-A. This naturally reduces the uncertainty on the base power exchange as any point within the ellipsoid must be feasible for all realizations of the uncertainty, meaning a larger ellipsoid would reserve more DERs flexibility. This leads to the following set of possible baseload powers, as in (6).
@@
    $p^{b}=(B_{1}^{b}D^{{b}^{-1}}(E^{b}\xi+(e^{b}-b(\zeta)))+B_{2}^{b}(K^{b}\xi+\gamma^{b}+L\zeta)$
  • For the controllable baseload case, two new variables $p_{0}^{b}$ and $y^{b}$ are introduced. These represent the expected slack power and controllable injections (16). In this case only the expected value $p^{b}_{0}$ is optimized.
@@ II-D Probabilistic Constraint Satisfaction
The available power flexibility in ADNs depends on the stochastic prosumption, as it influences both the network constraints and the capabilities of DERs. Failing to account for these uncertainties leads to an overestimation of the ADN’s power flexibility, while robustly ensuring the flexibility capacity is available can be overly conservative. Certain grid operators have already recognized this issue [16]. To allow DERs with uncertain capabilities to participate in the AS markets, the grid operator allows actors to offer flexibility services as long as the offered capacity is available at least 90% of the time [11] (i.e. according to the P90 requirement). To this end, the stochastic prosumption is handled in a probabilistic way in this work, with the availability requirement interpreted as a joint chance constraint (JCC) enforcing a 90% probability that the the full flexibility is available at each time step. As JCCs are generally hard to solve, we determine a robust uncertainty set for the uncertainty drivers that guarantees a certain coverage $1-\epsilon$ of the empirical probability distribution. Note that uncertainty sets can be obtained for each time step separately, considering as stochastic variables $\zeta_{t}^{u}$ (e.g. with $\zeta_{t}^{u}||\leq 1$ at time $t$ as was done in [9]). However, to guarantee a coverage of the joint probability distribution and jointly satisfy the problem constraints, a single uncertainty set for $\zeta=[\zeta^{u1},\zeta^{u1}...\>\zeta^{u|\mathcal{U}|}]$ should be considered for all uncertainty drivers over the full horizon. In this work, we assume a single uncertainty driver $\zeta^{u}$ suffices to model the stochastic fluctuations of all individual injections of a certain type (e.g. all PV injections are modelled with the same uncertain irradiance since, for distribution networks it is reasonable to assume they are located in the same geographical area.). Stochastic injections can be formulated as $p^{u}=\bar{p}^{u}+M\zeta^{u}\>\forall u\in\mathcal{U}$, with $M^{u}$ a linear mapping from the uncertainty drivers to the stochastic power injections [^4] Consider for example PV plants. The power can be approximated as $P^{cap}*$GHI, where the GHI uncertain. Ellipsoidal uncertainty sets are frequently used due to their simple parametric and numerical representation. Such uncertainty sets do not enforce simultaneous satisfaction of the worst case realization for all considered random inputs, thus often leading to less conservative solutions [1]. Furthermore, stochastic uncertainty often naturally allows an ellipsoidal representation, (e.g. uncertainty drivers with Gaussian distributions allow exact ellipsoidal uncertainty sets [3].) A first approach to determine an ellipsoidal uncertainty set is to fit a multivariate Gaussian distribution to the uncertain prosumption data. Ellipsoidal uncertainty sets can also be constructed to ensure a $1-\epsilon$ coverage of the empirical probability distribution of the available data. A disadvantage of ellipsoidal sets is that they may inflate the uncertainty set in certain directions, especially when extreme values for the different variables are correlated, such as PV at subsequent time steps. Therefore, we also consider a hyperbox uncertainty set covering the same proportion of the empirical samples. A comparison of these approaches is presented in Section III-A.
II-E Cost-Effective Flexibility Maximisation
Existing literature typically maximises the volume of the aggregation set when considering multiple time steps. However, when considering multiple ASs, the value of the services at each time step becomes relevant. Given that any $\xi||\leq 1$ can be selected, the realized benefit is not known a priori. Additionally, optimizing for the best case as an example would lead to the maximization of a norm, making the problem non-convex. Finally, as any point on the ellipsoid may be selected, the objective should reflect both the value of all points on the ellipsoid and the range within which the system operator can select the desired power reservation. To quantify this value, the different service prices are used. This reflects the willingness of system operators to pay and thus the service value for the system. Therefore the objective in (17) is selected as a proxy to maximize the flexibility value. Here,the quantity of flexibility offered at each time step is represented by the diagonal elements of $E^{s}$. Denoting the service prices at time steps $t$ as $\pi^{s}_{t}$, the objective becomes:
$J=\sum_{s}J^{s}\;=\sum_{s}\sum_{t}\pi^{s}_{t}E^{s}(t,t)$ (17)
In principle, the aggregation set may be modelled differently than through an ellipsoid. Two common approaches in the literature rely on the use of (i) polygonal aggregation or (ii) hyperboxes. However, in the case of a polygon, for which the parameters are unknown, the flexibility maximization problem becomes non-convex [18]. In the case of a hyperbox, the problem can be reformulated in a convex way but, for the considered objective, the aggregation set does not capture the tradeoff between different bids as a hyperbox enforces all bids to be satisfied simultaneously.
@@ II-E Cost-Effective Flexibility Maximisation
A sufficient condition for the convexity of the problem is that the costs for the DERs must be linear with respect to the allocated power capacity. This is a reasonable assumption as can easily be seen for the example of BESSs with equivalent cycles. For linear cost functions, the flexibility cost becomes:
@@
Finally, to avoid biasing the energy available in the baseload case and thus unfairly increasing the profit by emptying storage assets, two constraints are added.
@@
$\sum_{t}(c^{0,t}p^{b}_{0,t}-\sum_{r}c^{r^{T}}_{t}|p^{b,t}|)\geq 0$ (25)
The first one (24) enforces the integral of the difference of baseload power with respect to the expected net load to be zero. The second (25) ensures the power balancing accounts for operational costs, where $r$ indexes the different DERs. Including these cost-effectiveness considerations, the final problem is given by Problem (26). The impact of this controlled baseload on the problem constraints is modelled through (26m) and the influence of the uncontrollable prosumption and corresponding adjustment of the base power is included in (26g). Note that the subscript $u$ refers to the number of uncertainty sets to ensure the formulation remains generic. The influence of all the services and the selected baseload control is combined in (26h) and (26i) where $e^{0}$ and $\gamma$ disappear for the baseload case. We introduce auxiliary variables $\alpha^{i}$ for each constraint and $\alpha^{i}_{s}$ to constrain the impact of the different services on each constraint. Reformulating the problem with robust constraints and restricting the aggregated flexibility set to (quadrant) ellipsoids for all services yields the multiservice aggregation problem (26).
$\max_{E^{s},K^{s}}\quad\sum_{s}\sum_{t}\pi^{s}_{t}E^{s}(t,t)$ (26a)
@@
$(w^{0}_{2,i}L^{0}-\theta_{u,i})\zeta\leq\lambda_{i}\quad\forall\zeta\in\mathcal{U}$ (26g)
@@
$\sum_{t}(c^{0}_{t}p^{b}_{0,t}-\sum_{r}c^{r^{T}}_{t}|p^{b,t}|)\geq 0$ (26o)
Finally, depending on the representation of the uncertainties (26g), the effect of uncertainty is reformulated in (27a) and (27b) and (27b) for ellipsoid and hyperbox uncertainty sets respectively.
[…]
$\lambda\geq\sum_{u}(w^{0}_{2,i,u}L^{0}_{u}-\theta_{u,i})\zeta_{u}^{\max}$ (27b)
$\lambda\geq\sum_{u}(w^{0}_{2,i,u}L^{0}_{u}-\theta_{u,i})\zeta_{u}^{\min}$ (27c)
@@ III Results
This section demonstrates the approach for a number of test cases. First results for a the IEEE33 benchmark system are shown when considering only BESSs for the baseload and self-dispatching cases. Then the approach is extended to multiple feeders, interconnected to a common primary substation, showing its effectiveness when aggregating the flexibility of realistic distribution systems. Next, the problem is extended to different DERs, showing how they participate in the aggregation. Finally the impact of network constraints is discussed. Common inputs for the service prices and the uncertainty sets are used. Input data is obtained from the Réseau de Transports d’Electricité [39] (for the prices), Deutscher Wetterdienst [27] (for the PV production) and historical residential data from [25] for the loads. Figure 2 shows the inputs used for the case studies.
TABLE I: System Component Parameters
PV BESS
Node 5 11 14 18 20 25 28 32 8 17 25 33
Rating [kW] 100 200 100 500 100 50 200 150 1000 4000 2000 1000
Load
Node 2 3 4 6 8 9 10 11 13 15 16
Rating [kW] 90 120 60 200 60 60 45 60 120 60 60
Node 17 19 21 22 23 24 26 27 29 30 31
Rating [kW] 90 90 90 90 420 420 60 60 200 150 60
TABLE II: Power Rating of DERs in the IEEE33 network.
The results presented below all consider as flexibility services the provision of FCR and aFRR (up and down) as examples of symmetric and asymmetric services. To construct the uncertainty sets $1200$ from the $40x40=1600$ scenarios are used (i.e. combining the load scenarios with the weather ones) and $400$ are kept for a posteriori simulation of the flexibility provision.
@@
Figure 3: Uncertainty sets for the different ellipsoidal approaches and the hyperbox approach.
III-A Comparison of the Uncertainty Sets
This section shows the impact of the uncertainty representation on the obtained aggregated flexibility. Ellipsoidal and hyperbox uncertainty sets are obtained by ensuring a proportion $1-\epsilon$ (here set to $0.9$) of the in-sample scenarios are within the uncertainty sets. One approach to obtain such a set is based on fitting a multivariate Gaussian distribution to the prediction scenarios. As the load and PV uncertainties are not truly normally distributed, this leads to an overly conservative uncertainty set. Alternatively, robust ellipsoidal sets can be obtained by determining the minimum volume ellipsoid containing all points in prediction set [5]. Assuming this set is representative for the future realizations allows to obtain probabilistic uncertainty sets containing a proportion $1-\epsilon$ of all the points. For example, by computing the mutual distance between all points and selecting the one with the lowest maximum distance to the $1-\epsilon$ points with the smallest distance to this point. Figure 3 shows the boundaries of the ellipsoidal sets obtained in different ways. Representing the ellipsoids is not possible due to the large number of dimensions. Instead, the figure shows the range of possible values at all time steps for the different approaches, (i.e. the extreme points in all dimensions). The bounds for a robust ellipsoid considering all the in-sample scenarios are shown in yellow, the probabilistic uncertainty set assuming a Gaussian distribution is shown in blue and the set ensuring $90\%$ coverage of the in-sample scenarios is shown in green. The range of prediction scenarios and historical realizations used to obtain the uncertainty sets are also shown. Finally, the figure also shows the bounds of the hyperbox uncertainty set, which corresponds to the maximum and minimum values of the in-sample scenarios. Figure 4 shows the aggregated flexibility for all cases (only aFRR up is selected by the optimization problem in this case). Clearly, the hyperbox uncertainty set is the least conservative. The constraint violation rates in Table III confirm this trend, with all ellipsoidal sets leading to much lower violation rates than the allowed $10\%$.
Figure 4: Impact of uncertainty set on aggregated flexibility.
Aggregation Set Hyperbox 90% ellipsoid Robust Ellipsoid Gaussian Ellipsoid
Violation Rate [%] 6.25 0 0 0.75
TABLE III: Constraint Violation Rates